# A recursive approach for the determination of the nice sections of width three having a 4-crown stack as retract

### v1

@@ Abstract
The characterization of the finite minimal automorphic posets of width three is still an open problem. Niederle has shown that this task can be reduced to the characterization of the nice sections of width three having a non-trivial tower of nice sections as retract. In our article, we characterize those of them which have a 4-crown stack as retract and we develop a recursive approach for their determination. We apply the approach on a sub-class $\mathfrak{N}_{2}$ of nice sections of width three and determine all posets in $\mathfrak{N}_{2}$ with height up to six having a 4-crown stack as retract. For each integer $n\geq 2$, the class $\mathfrak{N}_{2}$ contains $2^{n-2}$ different isomorphism types of posets of height $n$.
@@ 1 Introduction
However, the description or characterization of minimal automorphic posets is still an open field. (Unless otherwise stated, “poset” always means “finite poset” in this article.) Since the pioneering work of Rival [7, 8], we know that a poset of height one does not have the fixed point property iff it contains a crown, because a crown of minimal length will be a retract of it. Brualdi and Da Silva [1] generalized this result by showing that a poset does not have the fixed point property iff it has a “generalized crown” as retract. With respect to posets of small width, Fofanova and Rutkowski [4] showed that a poset of width two does not have the fixed point property iff it has a 4-crown stack as retract. (The result extends even to chain-complete infinite posets.) Niederle [5] characterized in 1989 posets of width three not having the fixed point property by having a non-trivial retract being a “tower of nice sections”. Later on, in 2008, he extended his result to posets of width four [6].
@@
Unfortunately, it is quite difficult to break down Niederle’s abstract characterization of the fixed point property to concrete posets or classes of posets of width three. According to our knowledge the only real success in this field is a result of Farley [3] from 1997, who showed that the ranked nice sections of width three are exactly the 6-crown stacks, and that a 6-crown stack is minimal automorphic iff its height $h$ is not a multiple of three. Otherwise, it has a 4-crown stack of height $\frac{2}{3}h$ as retract. Schröder [10, Lemma 2.6] generalized one direction of this result by showing that a $2n$-crown stack of height $nm$ has a 4-crown stack of height $2m$ as retract. Moreover, in the case of $n\geq 3$, a $2n$-crown stack of height $h$ is not minimal automorphic if $h$ and $n$ have a non-trivial common factor [10, Prop. 2.7].
Due to the results of Niederle [5], the investigation of the minimal automorphic posets of width three can be restricted to the investigation of the nice sections of width three. In our article, we characterize those of them which have a 4-crown stack as retract and we develop and apply a recursive approach for their determination.
@@
After the preparatory Section 2, we deal with the classes $\mathfrak{N}$ of nice sections and $\mathfrak{C}$ of crowned sections in Section 3; $\mathfrak{N}$ is a proper subclass of $\mathfrak{C}$. We show in Theorem 3.4 that a poset $P\in\mathfrak{C}$ has a 4-crown stack as retract iff it can be decomposed by a retractive split in a certain way. This characterization gives raise to a recursive approach for the determination of the nice sections having a 4-crown stack as retract; it is described at the end of Section 3.
@@
In order to test the approach, we introduce the class $\mathfrak{N}_{2}$ of nice sections with horizon two in Section 4. For each integer $n\geq 2$, this class contains $2^{n-2}$ isomorphism types of posets of height $n$; the isomorphism types are described in Theorem 4.2. We characterize in Theorem 4.5 the retractive splits of posets $P\in\mathfrak{N}_{2}$ fulfilling the requirements of Theorem 3.4, and using this result, we determine in Section 5 by means of the recursive approach all posets $P\in\mathfrak{N}_{2}$ of height up to six having a 4-crown stack as retract.
@@
The characterization of the minimal automorphic posets of width three is regarded as a quite difficult task [9, p. 98], in parts even as “intractable” [3, p. 126]. Also the sparse sequence of publications dealing with the topic indicates the difficulties. Nevertheless, it gives some confidence that the nice section of width three used by Schröder [9, p. 99] for the illustration of the difficulties and considered as “nasty” by Farley [3, p. 136] is handled without a problem by the machinery developed in this article. It is the poset coded as 1011 in Section 5 and shown in Figure 7.
@@ 2 Preparation
$P(k,\ell)$ $:=P(k)\cup P(\ell),$
$P(k\rightarrow\ell)$ $:=\cup_{i=k}^{\ell}P(i).$
For $k_{1}\leq k_{2}$ being level indices of $P$, we call the sub-poset $P(k_{1}\rightarrow k_{2})$ a segment of $P$, a lower segment if $k_{1}=0$, and an upper segment if $k_{2}=h_{P}$.
For $n\in\mathbb{N}\setminus\{1\}$, a $2n$-crown is a poset with $2n$ points $x_{i,j}$, $i\in[0,n-1]$, $j\in\{0,1\}$, in which $x_{0,0}<x_{0,1}>x_{1,0}<\cdots>x_{n-1,0}<x_{n-1,1}>x_{0,0}$ are the only comparabilities between different points. We call a poset $P$ of height $\geq 1$ a $2n$-crown stack iff $P(k,k+1)$ is a $2n$-crown for all $k\in[0,h_{P}-1]$ and all other comparabilites are induced by transitivity.
@@
Given disjoint posets $P=(X,\leq_{P})$ and $Q=(Y,\leq_{Q})$, their ordinal sum is the poset $P\oplus Q:=(X\cup Y,\leq_{P}\cup\leq_{Q}\cup\;(X\times Y))$. For $P$ and $Q$ being antichains of size 3, we simply write “3333” as shortcut for the isomorphism type of $P\oplus Q$. Additionally, we write $3C$ for a poset consisting of three disjoint chains of height 1 without any additional comparabilities between different points. The shortcuts are illustrated in Figure 1.
For a mapping $f:X\rightarrow Y$, we denote by $f|_{Z}$ the pre-restriction of $f$ to a subset $Z$ of its domain and by $f|^{Z}$ its post-restriction to a subset $Z$ of its codomain with $f[X]\subseteq Z$. An order homomorphism $r:P\rightarrow P$ is called a retraction of the poset $P$ if $r$ is idempotent, and an induced sub-poset $R$ of $P$ is called a retract of $P$ if a retraction $r:P\rightarrow P$ exists with $r[X]=R$. For the sake of simplicity, we identify $r$ with its post-restriction and write $r:P\rightarrow R$. A poset $P$ has the fixed point property iff every retract of $P$ has the fixed point property [9, Th. 4.8]. We need the following well-known result about the existence of spanning retracts:
@@
Let $P$ be be a finite poset and $r:P\rightarrow R$ a retraction. There exists a retraction $v:P\rightarrow V$ with $V\simeq R$, $\min V\subseteq\min P$ and $\max V\subseteq\max P$.
@@
3 Nice sections and crowned sections
@@ 3.1 Definition and basic properties
  • The mapping $c_{k,j}\mapsto c_{k,(j+1)\!\!\!\mod 3}\;$ for all $k\in[0,h_{P}]$, $j\in\{0,1,2\}$ is an automorphism of $P$;
@@
For a section $P$ of width 3, the first and second property imply $c_{0,j}\lessdot\ldots\lessdot c_{h_{P},j}$ for all $j\in\{0,1,2\}$. We call the chains $C_{0},C_{1},C_{2}$ the main chains of $P$. Clearly, $P(k)=\{c_{k,0},c_{k,1},c_{k,2}\}$ for all $k\in[0,h_{P}]$. We define mappings $\lambda:P\rightarrow[0,h_{P}]$ and $\gamma:P\rightarrow\{0,1,2\}$ by setting $\lambda(c_{k,j}):=k$ and $\gamma(c_{k,j}):=j$, hence $\{p\}=P(\lambda(p))\cap C_{\gamma(p)}$ for all $p\in P$. Two consecutive levels of $P$ form a 6-crown or are of type $3C$. In what follows, the level-index function $\lambda$ and the chain-index function $\gamma$ always refer to the levels and main chains of $P$; for a retract $R$ of $P$, we thus have $\{x\}=P(\lambda(x))\cap C_{\gamma(x)}$ for all $x\in R$.
Niederle [5, p. 121 and p. 125] has shown that the decomposition of a tower of sections into sections is unique, and that a poset of width at most three has not the fixed point property iff it has a tower of nice sections as retract. We conclude that an automorphic poset $P$ of width at most 3 not being a tower of nice sections is not minimal automorphic. Moreover, for $P=S_{1}\oplus\cdots\oplus S_{n}$ being a tower of nice sections $S_{1},\ldots S_{n}$, it is not hard to see that $P$ is minimal automorphic iff $S_{i}$ is minimal automorphic for every $i\in[1,n]$. The objects we have to deal with in investigating minimal automorphic posets of width three are thus the nice sections of width three.
@@
Our first goal is the characterization of nice sections of width three having a 4-crown stack as retract. However, our results will concern a larger class of sections containing the nice sections of width three as a proper subclass:
@@
3.2 Four-crown stacks as retracts of crowned sections

### v2

# A recursive approach for the determination of the nice sections of width three which have a 4-crown stack as retract Version 8-3

@@ Abstract
The characterization of the finite minimal automorphic posets of width three is still an open problem. Niederle has shown that this task can be reduced to the characterization of the nice sections of width three having a non-trivial tower of nice sections as retract. In our article, we develop a recursive approach for the determination of those of them which have a 4-crown stack as retract. We apply the approach on a sub-class $\mathfrak{N}_{2}$ of nice sections of width three and determine all posets in $\mathfrak{N}_{2}$ with height up to six having a 4-crown stack as retract. For each integer $n\geq 2$, the class $\mathfrak{N}_{2}$ contains $2^{n-2}$ different isomorphism types of posets of height $n$.
@@ 1 Introduction
However, the description or characterization of minimal automorphic posets is still an open field. (Unless otherwise stated, “poset” always means “finite poset” in this article.) Since the pioneering work of Rival [7, 8], we know that a poset of height one does not have the fixed point property iff it contains a crown, because a crown of minimal length will be a retract of it. Brualdi and Dias Da Silva [1] generalized this result by showing that a poset does not have the fixed point property iff it has a “generalized crown” as retract. With respect to posets of small width, Fofanova and Rutkowski [4] showed that a poset of width two does not have the fixed point property iff it has a 4-crown stack as retract. (The result extends even to chain-complete infinite posets.) Niederle [5] characterized in 1989 posets of width three not having the fixed point property by having a non-trivial retract being a “tower of nice sections”. Later on, in 2008, he extended his result to posets of width four [6].
@@
Unfortunately, it is quite difficult to break down Niederle’s abstract characterization to concrete posets or classes of posets. As far as we know, the only real success in this field is a result of Farley [3] from 1997, who showed that the ranked nice sections of width three are exactly the 6-crown stacks, and that a 6-crown stack is minimal automorphic iff its height $h$ is not a multiple of three. Otherwise, it has a 4-crown stack of height $\frac{2}{3}h$ as retract. Schröder [10, Lemma 2.6] generalized one direction of this result by showing that a $2n$-crown stack of height $nm$ has a 4-crown stack of height $2m$ as retract. Moreover, in the case of $n\geq 3$, a $2n$-crown stack of height $h$ is not minimal automorphic if $h$ and $n$ have a non-trivial common factor [10, Prop. 2.7].
@@
Due to the results of Niederle [5], the study of the minimal automorphic posets of width three can be restricted to the investigation of the nice sections of width three. In our article, we develop and apply a recursive approach for their determination. After the preparatory Section 2, we deal with the classes $\mathfrak{N}$ of nice sections and $\mathfrak{C}$ of crowned sections in Section 3; $\mathfrak{N}$ is a proper subclass of $\mathfrak{C}$. We see that a poset $P\in\mathfrak{C}$ has a 4-crown stack as retract iff it has such a retract containing a split level, and we prove several results about the existence of such levels. We see in Lemma 3.11 that a split level of $R$ corresponds in $P$ to a certain decomposition by a retractive split, and based on this observation we develop at the end of Section 3 a recursive approach for the determination of the nice sections having a 4-crown stack as retract.
@@
In order to test the approach, we introduce the class $\mathfrak{N}_{2}$ of nice sections with horizon two in Section 4. For each integer $n\geq 2$, this class contains $2^{n-2}$ isomorphism types of posets of height $n$; the isomorphism types are described in Theorem 4.2. We characterize in Theorem 4.5 the retractive splits of posets $P\in\mathfrak{N}_{2}$ yielding a 4-crown stack as retract. Using this result, we determine in Section 5 by means of the recursive approach all posets $P\in\mathfrak{N}_{2}$ of height up to six which have a 4-crown stack as retract.
@@
The characterization of the minimal automorphic posets of width three is regarded as a difficult task [9, p. 98], in parts even as “intractable” [3, p. 126]. Also the sparse sequence of publications indicates these difficulties. Nevertheless, it gives some confidence that the nice section of width three used by Schröder [9, p. 99] for the illustration of the problems and considered as “nasty” by Farley [3, p. 136] is handled without a problem by the machinery developed in this article. It is the poset coded as 1011 in Section 5 and shown in Figure 9.
@@ 2 Preparation
$P(k,\ell):=P(k)\cup P(\ell)\quad\>\textrm{and}\>\quad P(k\rightarrow\ell):=\cup_{i=k}^{\ell}P(i).$
For level indices $k_{1}\leq k_{2}$ of $P$, we call the sub-poset $P(k_{1}\rightarrow k_{2})$ a segment of $P$, a lower segment if $k_{1}=0$, and an upper segment if $k_{2}=h_{P}$.
For $n\in\mathbb{N}\setminus\{1\}$, a $2n$-crown is a poset with $2n$ points $x_{i,j}$, $i\in[0,n-1]$, $j\in\{0,1\}$, in which $x_{0,0}<x_{0,1}>x_{1,0}<\cdots>x_{n-1,0}<x_{n-1,1}>x_{0,0}$ are the only comparabilities between different points. We call a poset $P$ of height at least 1 a $2n$-crown stack if $P(k,k+1)$ is a $2n$-crown for all $k\in[0,h_{P}-1]$ and all other comparabilites are induced by transitivity.
@@
Given disjoint posets $P=(X,\leq_{P})$ and $Q=(Y,\leq_{Q})$, their ordinal sum is the poset $P\oplus Q:=(X\cup Y,\leq_{P}\cup\leq_{Q}\cup\;(X\times Y))$. For $P$ and $Q$ being disjoint antichains of size 2 or 3, we simply write “#P#Q\#P\#Q” as shortcut for the isomorphism type of $P\oplus Q$, e.g., $22$ indicates a 4-crown. Additionally, we write $2C$ and $3C$ for a poset consisting of two (three) disjoint chains of height 1 without any additional comparabilities between different points. The shortcuts are illustrated in Figure 1.
For a mapping $f:X\rightarrow Y$, we denote by $f|_{Z}$ the pre-restriction of $f$ to a subset $Z$ of its domain and by $f|^{Z}$ its post-restriction to a subset $Z$ of its codomain with $f[X]\subseteq Z$. For $X^{\prime}\supseteq X$, we call a mapping $f^{\prime}:X^{\prime}\rightarrow Y$ an extension of $f$, if $f^{\prime}|_{X}=f$.
An order homomorphism $r:P\rightarrow P$ is called a retraction of the poset $P$ if $r$ is idempotent, and an induced sub-poset $R$ of $P$ is called a retract of $P$ if a retraction $r:P\rightarrow P$ exists with $r[X]=R$. For the sake of simplicity, we identify $r$ with its post-restriction and write $r:P\rightarrow R$. A poset $P$ has the fixed point property iff every retract of $P$ has the fixed point property [9, Th. 4.8]. We need the following well-known result about the existence of spanning retracts:
@@
Let $P$ be be a finite poset and $r:P\rightarrow R$ a retraction. There exists a retraction $v:P\rightarrow V$ with $V\simeq R$, $\min V\subseteq\min P$, $\max V\subseteq\max P$, and $V(1\rightarrow h_{V}-1)=R(1\rightarrow h_{R}-1)$.
@@
3 Nice sections
@@ 3.1 Definition and basic properties
  • The mapping $c_{k,j}\mapsto c_{k,(j+1)\!\!\!\mod 3}\;$ for all $k\in[0,h_{P}]$, $j\in\{0,1,2\}$, is an automorphism of $P$;
@@
For a section $P$ of width 3, the first and second property imply $c_{0,j}\lessdot\ldots\lessdot c_{h_{P},j}$ for all $j\in\{0,1,2\}$. We call the chains $C_{0},C_{1},C_{2}$ the main chains of $P$. Clearly, $P(k)=\{c_{k,0},c_{k,1},c_{k,2}\}$ for all $k\in[0,h_{P}]$. We define mappings $\lambda:P\rightarrow[0,h_{P}]$ and $\gamma:P\rightarrow\{0,1,2\}$ by setting $\lambda(c_{k,j}):=k$ and $\gamma(c_{k,j}):=j$, hence $\{p\}=P(\lambda(p))\cap C_{\gamma(p)}$ for all $p\in P$. Two consecutive levels of $P$ form a 6-crown or are of type $3C$. The level-index function $\lambda$ and the chain-index function $\gamma$ always refer to the levels and main chains of $P$; for a retract $R$ of $P$, we thus have $\{x\}=P(\lambda(x))\cap C_{\gamma(x)}$ for all $x\in R$.
Niederle [5, p. 121 and p. 125] has shown that the decomposition of a tower of sections into sections is unique, and that a poset of width at most three has not the fixed point property iff it has a proper retract being a tower of nice sections. We conclude that an automorphic poset $P$ of width at most 3 not being a tower of nice sections is not minimal automorphic. Moreover, if $P=S_{1}\oplus\cdots\oplus S_{n}$ is a tower of nice sections $S_{1},\ldots S_{n}$, it is not hard to see that $P$ is minimal automorphic iff $S_{i}$ is minimal automorphic for every $i\in[1,n]$. The objects we have to deal with in investigating minimal automorphic posets of width three are thus the nice sections of width three.
@@
3.2 Crowned sections and split levels
Our goal is a recursive approach for the determination of the nice sections of width three having a 4-crown stack as retract. However, our results will concern a larger class of sections containing the nice sections of width three as a proper subclass:
@@ 3.2 Crowned sections and split levels
For a retraction $r:P\rightarrow R$ of $P\in\mathfrak{C}$ onto a 4-crown stack, we are interested in proper lower or upper segments of $R$ which are the image of a proper lower or upper segment of $P$:
Definition 3.3.
Let $P\in\mathfrak{C}$, let $r:P\rightarrow R$ a retraction onto a 4-crown stack, and let $R(\ell)=\{a,b\}$ with $\lambda(a)\leq\lambda(b)$. We call $R(\ell)$ a down-split level if
$\ell<h_{R}\quad\>\textrm{and}\>\quad R(0\rightarrow\ell)=r[P(0\rightarrow\lambda(b))],$
and we call $R(\ell)$ an up-split level if
$0<\ell\quad\>\textrm{and}\>\quad R(\ell\rightarrow h_{R})=r[P(\lambda(a)\rightarrow h_{R})].$
In the retract of the poset on the left in Figure 2, the level set $R(0)$ is a down-split level, the level set $R(2)$ is an up-split level, and $R(1)$ is neither down- nor up-splitting. For the retract of the poset on the right of the figure, the down-split levels are $R(0)$ and $R(1)$ and the up-split levels are $R(1)$ and $R(2)$.
We agree on that for the rest of this section the poset $P$ is a crowned section and $r:P\rightarrow R$ is a retraction onto a 4-crown stack. This implies $h_{R}\geq 2$: Suppose to the contrary that $R$ is a 4-crown. For the corresponding spanning retraction $r^{\prime}:P\rightarrow R^{\prime}$, the poset $Q:=P(0,1)\cup P(h_{P})$ has the retraction $r^{\prime}|_{Q}:Q\rightarrow R^{\prime}$ onto the 4-crown $R^{\prime}$. But this is impossible because $Q$ is a 6-crown or a 6-crown stack of height 2, or it contains a point $x\in Q(1)$ with $R^{\prime}(0)<x<R^{\prime}(1)$.
In our investigation of split levels, the level sets $R(0)$, $R(1)$, $R(h_{R}-1)$, and $R(h_{R})$ frequently will play a special role. We notate the results for $R(0)$ and $R(1)$ only; the results for $R(h_{R})$ and $R(h_{R}-1)$ are dual.
We start with a simple observation: Because the down-sets $r^{-1}(a)$, $a\in R(0)$, are pairwise disjoint,
$R(0)\cap P(0)\not$$\emptyset$. (1)
Lemma 3.4.
Assume that the level set $R(\ell)$ is contained in a single level set of $P$. For $1\leq\ell\leq h_{R}-1$, the level set $R(\ell)$ is a split level, and for $\ell=0$, one of the level sets $R(0)$ and $R(1)$ is a split level.
Proof.
Let $R(\ell)$ be contained in $P(k)$ and let $z\in P(k)\setminus R(\ell)$.
  • $\ell\in[1,h_{R}-1]$: Then $k\in[1,h_{P}-1]$. Depending on $\lambda(r(z))\leq k$ or $\lambda(r(z))\geq k$, the set $R(\ell)$ is a down-split level or an up-split level.
  • $\ell=0$: Then $k=0$ due to (1). If $r(z)\in R(0)$, then $R(0)$ is a down-split level. And in the case of $r(z)\notin R(0)$, we have $R(0)<x$ or $z<x$ for all $x\in P(1\rightarrow h_{P})$ because $P$ is crowned, and $R(1)$ is an up-split level.
Lemma 2.1 and Lemma 3.4 together yield that a crowned section has a 4-crown stack as retract iff it has a 4-crown stack as retract containing a split level. On the one hand, this inconspicuous result brings our recursive approach into motion in Section 3.3. On the other hand, this approach becomes more effective and the concept of split levels more applicable if we can show that frequently there exist additional split levels not induced by the action of $r$ on $P(0)$ and $P(h_{P})$. In order to separate induced split levels from independent ones, we discriminate two cases:
  • $R(0)\subset P(0)$: Looking at the proof Lemma 3.4, we realize that if $R(0)$ is not a down-split level, then $R(1)$ is forced to be an up-split level.
  • $R(0)\not\subset P(0)$: Due to (1), there exists a point $a\in R(0)\cap P(0)$. For the point $b\in R(0)\setminus\{a\}$, we have $r[P(0)\setminus\{a\}]=\{b\}$, hence $r[P(1)\setminus\{c_{1,\gamma(b)}\}]\subseteq R(1\rightarrow h_{R})$. Nevertheless, two alternatives remain for the level set $R(1)$: it can be an up-split level or no split level at all.
In consequence, we define
Definition 3.5.
We call a split level $R(\ell)$ induced if
  • $\ell=0$ or $\ell=h_{R}$;
  • $\ell=1$, $R(0)\subset P(0)$, and $r[P(0)]\not\subset R(0)$;
  • $\ell=h_{R}-1$, $R(h_{R})\subset P(h_{P})$, and $r[P(h_{P})]\not\subset R(h_{R})$.
Otherwise, we call $R(\ell)$ independent.
For the rest of this section, we investigate structures which ensure the existence of independent split levels. We start with
Lemma 3.6.
Let $R(\ell)=\{a,b\}$ with $\lambda(a)<\lambda(b)$.
  1. 1.
    $\ell=0$: The level set $R(1)$ is an independent up-split level if
    $a$ $<P(\min\lambda[R(1)]).$
  2. 2.
    $1<\ell<h_{R}$: If
    $a$ $<P(\min\lambda[R(\ell+1)])$
    $\>\textrm{and}\>\quad b$ $>P(\max\lambda[R(\ell-1)]),$
    then $R(\ell-1)$ is a down-split level or $R(\ell+1)$ is an up-split level.
Proof.
1) We have $\lambda(a)=0$ according to (1). Due to $P(0)\setminus\{a\}<b$, we have $r(x)=b$ for all $x\in P(0)\setminus\{a\}$, thus $R(0)<r(y)$ for all $y\in P(\min\lambda[R(1)])$, and $R(1)$ is an up-split level. It is an independent one because of $R(0)\not\subset P(0)$.
2) Let $k:=\lambda(a)$, $k^{-}:=\max\lambda[R(\ell-1)]$, $j:=\lambda(b)$, and $j^{+}:=\min\lambda[R(\ell+1)]$. Let $x\in P(k)\setminus\{a\}$. If $r(x)\in\{b\}\cup R(\ell+1\rightarrow h_{R})$, then $P(k,j^{+})\simeq 33$ yields $R(\ell)<r(z)$ for all $z\in P(j^{+})$, and $R(\ell+1)$ is an up-split level.
Now assume $r(x)\in R(0\rightarrow\ell-1)\cup\{a\}$ for all $x\in P(k)\setminus\{a\}$. Then $r(z)\in R(0\rightarrow\ell-1)\cup\{a\}$ for all $z\in P(k^{-})$, and $P(k^{-},j)\simeq 33$ additionally yields $r(z)<b$, hence $r(z)\in R(0\rightarrow\ell-1)$, and $R(\ell-1)$ is a down-split level.
The retracts in Figure 2 illustrate the second part. In both, the level set $R(1)$ fulfills the assumptions, and $R(0)$ or $R(2)$ has to be a split level. In fact, both are.
Figure 3: The labeling of the points in $R(\ell,\ell+1)$ used in Lemma 3.7, Corollary 3.8, and Proposition 3.9. We assume $\lambda(a)\leq\lambda(b)$ and $\lambda(v)\leq\lambda(w)$.
For the rest of this section, we deal with a 4-crown $R(\ell,\ell+1)$. As indicated in Figure 3, we denote the points of $R(\ell)$ with $a$ and $b$ and assume $\lambda(a)\leq\lambda(b)$, and we denote the points of $R(\ell+1)$ by $v$ and $w$ and assume $\lambda(v)\leq\lambda(w)$.
Lemma 3.7.
Let $0\leq\ell\leq h_{R}-1$.
1) $P(\lambda(a),\lambda(w))\simeq 33$.
2) If $P(\lambda(b),\lambda(v))\simeq 33$, then $R(\ell)$ is a down-split level, $R(\ell+1)$ is an up-split level, and at least one of these split levels is independent.
3) If $P(\lambda(b),\lambda(v))$ is a 6-crown and $P(\lambda(b),\lambda(w))\simeq 33$, then $R(\ell)$ is a down-split level or $R(\ell+1)$ an up-split level.
Proof.
1) The points $a,c_{\lambda(a),\gamma(b)}$ and $w,c_{\lambda(w),\gamma(v)}$ form a 4-crown.
2) The assumption yields $P(\lambda(b))<R(\ell+1)$ and $R(\ell)<P(\lambda(v))$. In consequence, $R(\ell)$ is a down-split level and $R(\ell+1)$ an up-split level. For $1\leq\ell\leq h_{R}-2$, both split levels are independent, and for $\ell=0$, $R(0)$ being a down-split level makes $R(1)$ independent. For $\ell=h_{R}-1$, apply the dual argument.
3) The point $v$ has two lower covers $b$ and $z$ in $P(\lambda(b)),\lambda(v))$, and we have $z<R(\ell+1)$, hence $r(z)\in R(0\rightarrow\ell)$. For the point $y\in P(\lambda(b))\setminus\{b,z\}$, we have $y<P(\lambda((v))\setminus\{v\}$. If $r(y)\in R(0\rightarrow\ell)$, then $R(\ell)$ is a down-split level, and in the case of $r(y)\in R(\ell+1\rightarrow h_{R})$, the level $R(\ell+1)$ is an up-split level. The addendum follows directly from the definition of an independent split level.
Figure 4: Illustration of Corollary 3.8.
Corollary 3.8.
Let $1\leq\ell\leq h_{R}-2$ and let $P(\lambda(b),\lambda(v))$ form a 6-crown. Furthermore, as shown in Figure 4a, let $b^{\prime}$ be the point in $P(\lambda(b))\setminus\{b\}$ with $b^{\prime}<v$, and let correspondingly $v^{\prime}$ be the point in $P(\lambda(v))\setminus\{v\}$ with $v^{\prime}>b$.
If neither $R(\ell)$ nor $R(\ell+1)$ is splitting, then the poset induced by the points in $P(\lambda(a))\cup P(\lambda(b))\cup P(\lambda(v))\cup P(\lambda(w))$ is isomorphic to the poset shown in Figure 4b. Furthermore, $r(v^{\prime})=b$ and $r(b^{\prime})=v$.
Proof.
The relation $\lambda(a)<\lambda(b)$ is due to Lemma 3.4. Because of the dual of Lemma 3.7.3, the level pair $P(\lambda(a),\lambda(v))$ cannot be of type 33, and $P(\lambda(a),\lambda(v))$ must be a 6-crown. This implies $P(\lambda(a),\lambda(b))\simeq 3C$, and $a<v$ yields $a<b^{\prime}$. Due to Lemma 3.7.1, the type of $P(\lambda(a),\lambda(w)))$ is 33. Finally, $r(v^{\prime})\not=b$ would yield $r(v^{\prime})\in R(\ell+1\rightarrow h_{R})$, and depending on the image of the point in $P(\lambda(b))\setminus\{b,b^{\prime}\}$, the level set $R(\ell)$ would be a down-split level or $R(\ell+1)$ an up-split level. The rest is dual.
Proposition 3.9.
Let $R$ be a spanning retract and let $P(k,k+1)$ and $P(k+1,k+2)$ both be 6-crowns for a level index $k\in[1,h_{P}-3]$. With
[…]
the level set $R(\ell)$ is a down-split level or the level set $R(\ell+1)$ is an up-split level, and at least one of these split levels is independent.
Proof.
Because $R$ is spanning, the level index $\ell$ is well-defined with $\ell\leq h_{R}-1$. With our convention about the points in $R(\ell,\ell+1)=\{a,b,v,w\}$, the definition of $\ell$ implies $\lambda(a)\leq k$, $\lambda(b)\leq k+1$, $\lambda(v)\geq k+1$, and $P(\lambda(b),\lambda(v))\not\simeq 3C$.
Assume $\ell=0$. Then $\lambda(b)=0$ and $\lambda(v)\geq k+1\geq 2$ yield $P(\lambda(b),\lambda(v))\simeq 33$, and Lemma 3.7.2 delivers the result. The proof for $\ell+1=h_{R}$ is dual.
Now let $1\leq\ell\leq h_{R}-2$. In the case of $P(\lambda(b),\lambda(v))\simeq 33$, use Lemma 3.7.2 again.
Suppose thus that $P(\lambda(b),\lambda(v))$ is a 6-crown. If the type of $P(\lambda(a),\lambda(v))$ or $P(\lambda(b),\lambda(w))$ is 33, then apply Lemma 3.7.3 or its dual. Otherwise, we must have $\lambda(v)=k+1=\lambda(w)$. According to Lemma 3.4, $R(\ell+1)$ is an up-split level, hence independent, or a down-split level, which is independent for $\ell+2<h_{R}$.
Let thus $R(\ell+1)$ be a down-split level and $\ell+2=h_{R}$. Due to $R(\ell+1)\subset P(k+1)$, we have $R(h_{R})\not\subset P(k+2)$, thus $h_{P}\geq k+3$, and $R(\ell+1)<P(h_{P})$ follows. We conclude $r[P(h_{R})]=R(h_{R})$, and $R(\ell+1)$ is independent.
3.3 A recursive approach
[…]